OpenStudy (anonymous):

Choose all answers that are correct. 1. (1 pt) Which equations are true? Choose all answers that are correct. A. m∠3 = m∠1 + m∠2 B. m∠3 = m∠1 – m∠2 C. m∠1 = m∠2 + m∠3 D. m∠1 = m∠2 – m∠3

OpenStudy (anonymous):

only 2 answers

OpenStudy (anonymous):

a

OpenStudy (anonymous):

i need two answers

OpenStudy (anonymous):

a and ..

OpenStudy (anonymous):

so a. are b. are a are c.

OpenStudy (anonymous):

are a and d

OpenStudy (anonymous):

@TheCleverOne What do you think it is

OpenStudy (anonymous):

um hello ?

OpenStudy (anonymous):

A and D i think 0-0

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

im thinking d

OpenStudy (anonymous):

O-O THX

OpenStudy (anonymous):

Anytime:)

OpenStudy (anonymous):

Which equation can be used to find the value of h? A. 135 – 60 = h B. 135 + h = 60 C. 60 + 135 = h D. 60 – h = 135

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

im thinking not b,c, or d

OpenStudy (anonymous):

b,c, or d

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

Yes:)

OpenStudy (anonymous):

thx @TheCleverOne

OpenStudy (anonymous):

Welcome:)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

∠1 is decomposed into two nonoverlapping angles, ∠2 and ∠3. Let m∠1 = 130° and m∠3 = 75°. What type of angle is ∠2? A. acute B. obtuse C. right D. straight

OpenStudy (anonymous):

um

OpenStudy (anonymous):

im thinking right angle since 130 is obtuse and 75 is acute

OpenStudy (anonymous):

130 is obtuse and 75 is acute

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yes i said that

OpenStudy (anonymous):

c

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Welcome:) and i said the answer first:)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the missing angle measure? m∠1 = ° m∠2 = 75° m∠3 = 50° Three angles labeled 1, 2, and 3 are defined by rays. Angles 2 and 3 are adjacent. The non-shared rays from angles 2 and 2 form angle 1.

OpenStudy (anonymous):

the pic

OpenStudy (anonymous):

Angle 1 has to the biggest

OpenStudy (anonymous):

so a b c or d

OpenStudy (anonymous):

wait there are no choices

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then why did you say "a b c or d"

OpenStudy (anonymous):

What is the missing angle measure? m∠1 = ° m∠2 = 75° m∠3 = 50° Three angles labeled 1, 2, and 3 are defined by rays. Angles 2 and 3 are adjacent. The non-shared rays from angles 2 and 2 form angle 1.

OpenStudy (anonymous):

I guess 100 because the angles go 50, 75 counting by 25 so 50, 75, 100

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

and 1 looksz twice the size of 2

OpenStudy (anonymous):

What is the missing angle measure? m∠1 = 172° m∠2 = 125° m∠3 = ° Three angles labeled 1, 2, and 3 are defined by rays. Angles 2 and 3 are adjacent. The non-shared rays from angles 2 and 2 form angle 1.

OpenStudy (anonymous):

Anytime:)

OpenStudy (anonymous):

What is the missing angle measure? m∠1 = 172° m∠2 = 125° m∠3 = ° Three angles labeled 1, 2, and 3 are defined by rays. Angles 2 and 3 are adjacent. The non-shared rays from angles 2 and 2 form angle 1.

OpenStudy (anonymous):

pic?

OpenStudy (anonymous):

help pls

OpenStudy (anonymous):

What is the missing angle measure? m∠1 = 172° m∠2 = 125° m∠3 = ° Three angles labeled 1, 2, and 3 are defined by rays. Angles 2 and 3 are adjacent. The non-shared rays from angles 2 and 2 form angle 1.

OpenStudy (anonymous):

is there a pic

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

up top in blue

OpenStudy (anonymous):

So 78

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

Which equations are true? Choose all answers that are correct. A. m∠2 + m∠1 = m∠3 B. m∠2 + m∠3 = m∠1 C. m∠3 – m∠2 = m∠1 D. m∠1 – m∠2 = m∠3

OpenStudy (anonymous):

the pic

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Which equations are true? Choose all answers that are correct. A. m∠2 + m∠1 = m∠3 B. m∠2 + m∠3 = m∠1 C. m∠3 – m∠2 = m∠1 D. m∠1 – m∠2 = m∠3